Ten years ago, Jan was four times as old as Michael. Now, he is only twice as old as Michael. How old is Jan?
Answer by TimothyLamb(4379)

j - 10 = 4(m - 10)
j = 2m
---
j - 10 = 4(m - 10)
2m - 10 = 4m - 40
2m = 30
---
m = 15
j = 30
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php